Please help me solve this problem ((( X2 = 25))) (the 2 is a little 2 like 2nd power
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

X^2 = 25
x^2-25=0
(x-5)(x+5)=0
x=5 or x=-5
Cheers,
stan H.

Answer by checkley71(8403)

X~2 MEANS X TIMES X OR IN YOR CASE X TIMES X = 25 NOW WHAT 2 NUMBERS MULTIPLIED TOGETHER GIVE YOU 25? ANSWER 5*5=25 THUS 5~2=25